A rectangular garden is 20 ft longer than it is wide. Its area is 2925 . What are its dimensions?
Found 2 solutions by solver91311, checkley77:
Answer by solver91311(24713)



I already answered this one -- just a few minutes ago, in fact. Sorry, but the answer hasn't changed.

Let represent the width. Then represents the length. Since the area of a rectangle is the length times the width:



Put the quadratic in standard form:



Now all you need to do is solve the quadratic and take the positive root as the width and add 20 to get the length. Hint: It factors.

-45 times 65 = -2925


John


Answer by checkley77(12844)

W=L+20
LW=2925
L(L+20)=2925
L^2+20L-2925=0
(L-45)(L+65)=0
L-45=0
L=45 FT. IS THE LENGTH.
W=45+20=65 FT. IS THE WIDTH.
PROOF:
45*65=2925
2925=2925